OpenStudy (anonymous):

who can help me find the x intercepts to this problem: 2x2 + 10x = –12

OpenStudy (anonymous):

\[2x^{2}+10x+12=0\] \[2(x^{2}+5x+6)=0\] Divide both sides by two. \[(x+2)(x+3)=0\] So the answer is \[x=-2, -3\]

OpenStudy (anonymous):

ohh ok i think i understand now thx